OpenStudy (anonymous):

integration of 2sec2t

OpenStudy (callisto):

\[\int2sec2t dt\]\[=\int(2sec2t)\times\frac{sec2t+tan2t}{sec2t+tan2t}dt\]\[=\int2\frac{sec^22t+(sec2t)tan2t}{sec2t+tan2t}dt\]\[=\int\frac{1}{sec2t+tan2t}d(sec2t+tan2t)\]\[=...\]

OpenStudy (amistre64):

sec^2(t) or sec(2t) ?